OpenStudy (abdela25):

HELP!! quick question...GIVING OUT MEDALS! write an equation in slope-intercept form for the line that passes through the given points AND is PERPENDICULAR to the given line (5, -1); y=4x-7

OpenStudy (abdela25):

what?

OpenStudy (anonymous):

Do you need work?

OpenStudy (abdela25):

yeah i need some help and a quick explanation please

OpenStudy (anonymous):

ok. I will draw it out.

OpenStudy (abdela25):

ok would be appreciated if it was also explained in words too

OpenStudy (anonymous):



OpenStudy (anonymous):

Sorry about the typo

OpenStudy (abdela25):

yeah no probelm

OpenStudy (abdela25):

but one question, how is this perpendicular to the given line

OpenStudy (anonymous):

ok what you do is you take the slope of the original line and when you do the negative reciprocal of it, it makes the slope perpendicular.

OpenStudy (abdela25):

the given line was the points i listed and i dont understand how this is perpendicular to these points

OpenStudy (anonymous):

It is trust me. go ahead and graph it you will find out

OpenStudy (abdela25):

but isnt that what you do if your trying to find a parallel line to the given points

OpenStudy (anonymous):

no parallel line is the same slope different b point.

OpenStudy (abdela25):

oh okay i get it

OpenStudy (abdela25):

could you help me with another question and also check some of my before answers ?

OpenStudy (anonymous):

uh sure. IS this question closed?

OpenStudy (abdela25):

no not yet its still open and i rewarded you a medal

OpenStudy (anonymous):

k great. What is your next question?

OpenStudy (abdela25):

one quick question before we proceed, for this problem you listed b=-1+5/4 and then said find what B equals, is that what b would equal or how do i find b

OpenStudy (anonymous):

b will equal to -1+5/4. Whatever you get will equal to the b in the perpendicular equation

OpenStudy (abdela25):

oh okay cool

OpenStudy (abdela25):

another question about the perpendicular stuff

OpenStudy (abdela25):

same question different points

OpenStudy (abdela25):

write an equation in slope-intercept form for the line that passes through the given point AND is PERPENDICULAR to the given line (4, -2); y=3

OpenStudy (abdela25):

@elliecosme @nannerbug39 if you guys cant help with this please stick along so you can check answers for a couple questions after this one thanks

OpenStudy (anonymous):

ok. Think logically. To get a perpendicular line, you know the line must be x= to something. Now, it says 4,-2 so what you know that 4 must be x.

OpenStudy (abdela25):

and @Daredevil666 too

OpenStudy (anonymous):

ok ill watch what you post to help

OpenStudy (anonymous):

ok now the equation is y= 0 and x=4. so y= x+ b. So 0= 4+ b. b= -4. SO the equation should be x=4 + b

OpenStudy (abdela25):

how is y=0

OpenStudy (abdela25):

you say y= to 0 but i dont see a zero in the points at all could you please explain how y = 0 @Daredevil666

OpenStudy (anonymous):

y must be 0 because it has to be perpendicular to the y line

OpenStudy (abdela25):

oh okay i see but if y has to be 0 then why does it say (4, -2) in the given points -2 should be the y and the line should be perpendicular to these points correct?

OpenStudy (anonymous):

oh my bad. I am sorry. I made a mistake. y gives you where the line will be graph point 4, -2 and draw a line going to the right and left.

OpenStudy (anonymous):

That will give you the line which you are supposed to have. Your y intercept would be -2

OpenStudy (abdela25):

so it would look like this: y=-2 x=4 so y=x+b or -2=4+b and b=-4 correct?

OpenStudy (abdela25):

@nannerbug39 what do you think so far

OpenStudy (anonymous):

no b = -2. Your equation should be x=4 and y intercept = -2

OpenStudy (abdela25):

so what does y equal

OpenStudy (abdela25):

if b=-2 and x=4 what would y equal...still zero?

OpenStudy (abdela25):

@primeralph what do you think

OpenStudy (abdela25):

you seem like you know some math with all of those fans and medals @primeralph any help you can give

OpenStudy (anonymous):

i believe y should still equal o due to it having to be perpendicular to the y line

OpenStudy (abdela25):

oh okay so the only thing that changed was b? from -4 to -2 after the mistake was found?

OpenStudy (anonymous):

but if x-y+b=0